OpenStudy (anonymous):

If a line crosses the y-axis at (0, -3) and has a slope of 3, what is the equation of the line?

OpenStudy (apolloschariot):

so you can use point slope form for this one. do you know what it is?

OpenStudy (anonymous):

no,i dont know that

OpenStudy (apolloschariot):

y – y1 = m(x – x1)

OpenStudy (apolloschariot):

so your y-coordinate goes into the y1 and your x-coordinate goes into the x1, and m is where you plug in the slope

OpenStudy (apolloschariot):

so what does that look like?

OpenStudy (anonymous):

-3- -3=3(0-0) ?

OpenStudy (apolloschariot):

kind of. y+3=3(x-0)

OpenStudy (anonymous):

So the y and x stay the same but x1 and y 1 change?

OpenStudy (apolloschariot):

yes. you only plug in the coordinates for the x and y that have a 1 with it :)